write the class mark of the class 29-34
Answer by stanbon(75887)

write the class mark of the class 29-34
---------
The class mark is the middle value in the class.
----
class mark = (29+34)/2 = 63/2 = 31.5
===========================================
Cheers,
Stan H.
=============